OpenStudy (anonymous):

My calculatior only displays eight digits so I need help with this problem: x+x+500,000=625,000,000.

OpenStudy (anonymous):

312250000

OpenStudy (anonymous):

What is 312,250,000+500,000?

OpenStudy (anonymous):

Here's how to do this without a calculator. Using scientific notation, convert 500,000 and 625,000,000 to the following:\[500,000=5\times10^{5}\]\[625000000=625\times10^{6}\] Scientific notation means that if the exponent in the 10 is positive, you move the decimal point in the coefficient at the end of the number to the right that many times. For example 625 x 10^6 = 625. move the dot 6 places to the right filling in zeros as you go. After 3 dot moves: 625000. After 6 dot moves: 625000000. = 625 x 10^6 Now rewrite the equation as follows:\[x+x+5\times10^5=625\times10^6\] 1) add the x's to get 2x\[2x+5\times10^5=625\times10^6\] 2) subtract 5 x 10^5 from both sides of the equation\[2x=625\times10^6-5\times10^5\] 3) On the right side of the equal sign, factor out a 10^5 leaving you with\[2x=10^5(625\times10-5)=10^5(6245)\] 4) Divide both sides of the equation by 2\[x=10^5(6245)/2\] 5) Divide 6245 by 2 first to get\[x=10^5(3122.5)=3122.5\times10^5\] 6) Move the dot 5 places to the right After 3 moves: 3122500. After 2 more moves: 312250000. And there is your answer 312250000